# SLIFER: Investigating Performance and Robustness of Malware Detection Pipelines

### v1

@@ 3.4. SLIFER Implementation
We now present how we serialize the components we have described. SLIFER processes input programs by sequentially passing them from module to module, halting computations if one of those modules raises an alert. If one of the modules crashes due to pre-processing errors, SLIFER skips that module by passing the input to the next one. This can happen for EMBER-GBDT and Nebula, which require heavy pre-processing. We discuss classification error management in section Sect. 5, by showing that it is better to label samples as benign in case of pre-processing crashes.
@@
For the first YARA module, we collect 2.7k rules available in open GitHub repositories, accessed until November 2023333https://github.com/bartblaze/Yara-rules/tree/master/rules [^4] https://github.com/elastic/protections-artifacts/tree/main/yara/rules [^5] https://github.com/malpedia/signator-rules/tree/main/rules [^6] https://github.com/Neo23x0/signature-base/tree/master/yara [^7] https://github.com/Yara-Rules/rules/tree/master/malware. For the machine learning static analysis module, we leverage a library called secml_malware (Demetrio and Biggio, 2021) [^8] https://github.com/pralab/secml_malware which wraps both MalConv and GBDT-EMBER models. Instead of training both from scratch, we leverage pre-trained open-source implementations of both  [^9] https://github.com/endgameinc/malware_evasion_competition. As the last module of SLIFER leverages the Nebula model provided by its original repository [^10] https://github.com/dtrizna/nebula. Emulation, which is the core of Nebula, is achieved through Speakeasy (Mandiant. Speakeasy: portable, modular, binary emulator designed to emulate Windows kernel and user mode malware, 2022), a Windows kernel emulation library, which generates a behavioral report in a JSON file.
4. Experimental Setting
@@
Datasets. In our experiments, we employed two different datasets, that deferred in cardinality, malware families, benignware, and period of collections. The first one ($\mathcal{D}_{1}$) is composed of 5k malware, collected from VirusTotal before year 2018, and 3.5k benignware, harvested from GitHub and clean Windows installations. We show its composition in Table 1, drafted again utilising VirusTotal. As regards malware families, this dataset is unbalanced, with predominance of two specific families among all the others. While this composition might seem unsuitable for experiments, we believe it can be treated as a realistic snapshot of real production environments, where there is no control over the incoming threats to analyse.
@@
Table 1. Composition of $\mathcal{D}_{1}$ divided by malware class. There are 551 samples we were unable to correctly label.
@@
Table 2. Composition of $\mathcal{D}_{2}$ (Speakeasy dataset (Trizna, 2022)) by malware class.
The second dataset ($\mathcal{D}_{2}$) reflects the test set of dynamic analysis records published by Trizna (Trizna, 2022), comprising 10K benign-ware and 17.5K malware samples spanned across seven types (such as ransomware, trojans, keyloggers, etc.) as shown in Table 2. The dataset was collected in April 2022 by partnering with an undisclosed security vendor. We re-collected PE files from public data sources based on released hashes to perform static and YARA level analysis applicable only to raw PE bytes. Contrary to $\mathcal{D}_{2}$, this dataset has perfect balances between malware families, making it a good baseline for fairly assessing performances across different techniques.
@@
Mean Detection Time (MDT) to evaluate the mean time that each module/model needs to analyse an input sample and to make a decision. We compute this metric on a subset of our datasets.
Error Rate (ER) of each module/model to estimate the percentage of impossible-to-analyse samples. We also provide an experimental explanation of the best way to label those samples, since we must make a decision in a production scenario.
@@
Comparison between SLIFER and single models. In the first experiment, we test the predictive capabilities of SLIFER on our two datasets, comparing it against each separate module. In doing this we motivate the choice of our proposed architecture, especially for the dynamic analysis module. We show that the chosen Nebula is better than a competitor model, named Neurlux (Jindal et al., 2019), a LSTM model trained on dynamic reports extracted from emulation. Also, we improve the comparison by including Quo.Vadis (Trizna, 2022) which is a hybrid-analysis model that merges both static and dynamic features. On the contrary to SLIFER, Quo.Vadis introduces a monolithic structure with single forward- and back- propagation path, processing static and dynamic components simultaneously, further employing a ”meta-model” that accumulates representations from both analyses types to classify Windows executable samples.
@@
Effect of dynamic analysis in SLIFER Classification. Similarly to the analysis performed by Dambra et al. (Dambra et al., 2023), we investigate the efficacy of dynamic analysis. Differently from their setting, here we use emulation as last step of the analysis without concatenating together both static and dynamic features. Thereafter, we test the performances of SLIFER in two cases: in the former, we do not include the dynamic module, and in the latter, we test the full pipeline, including the model chosen after the analysis of the two separate models. Lastly, we also include SLIFER performance on detection divided per malware family contained in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
@@
Detection Time Analysis. We here want to compare the mean speed of each separate module or model, aiming to analyze the advantages and disadvantages in terms of time. We calculate the MDT metric on two subsets of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. The first one comprises 500 malware samples from $\mathcal{D}_{1}$ and 500 malware samples from $\mathcal{D}_{2}$. Similarly, we take 500 benignware samples from the two datasets. We also compute the standard deviation to show how the detection time can vary. For this evaluation, we separate malware and benignware samples because they differ in file dimension: benignware are usually bigger than malware, and this enlarges emulation, processing and classification times. Moreover, benign samples always pass through all SLIFER’s modules by design, and we prefer to evaluate times separately. To differ the two metrics, we denote with MDTm and MDTg the calculation of MDT for malware and benignware subsets respectively.
@@
Pipeline Robustness Evaluation. We test SLIFER in terms of robustness. We evaluate the capabilities of adversarial examples to deceive the pipeline. We decided to test a transfer attack approach, supposing to be an attacker knowing that the GBDT-EMBER model (which proves to be the best performing in most scenarios in terms of TPR and FPR, as we are going to discuss in Sect. 5) is used inside the pipeline. To do so, we rely on GAMMA, a state-of-the-art black-box attack (Demetrio et al., 2021a) based on section injection. This attack iteratively requests classification of perturbed samples to optimize the amount of injected content as sections, until either evasion is achieved or the query budget is terminated. We took 1k GBDT-classified malicious samples from $\mathcal{D}_{2}$ and injected 75 benign sections picked from $\mathcal{D}_{2}$ benignware samples. We set $\lambda=1\times 10^{-6}$ and we set the maximum number of queries to 100. These sections contain read-only data used by programs at runtime, but no further code is included inside adversarial EXEmples.
5. Experimental Results
We now describe all the findings we gathered from the experiments described in Sect. 4.
5.1. Single Modules Classification
@@
We test the performance of all the single modules and models with both $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ datasets in our possession, and we present results of each module in Table 3. We report the values of the described metrics when modules consider input as benignware in presence of pre-processing errors, by also including in brackets metrics computed considering classification errors as malware.
@@
Module TPR FPR ER
Signatures 0.15 5.2 ×$\times 10^{-3}$ 0
MalConv 0.78 2.0 ×$\times 10^{-2}$ 0
GBDT 0.91 (0.92) 4.1 ×$\times 10^{-3}$ (0.12) 5.4%
Nebula 0.23 (0.61) 2.3 ×$\times 10^{-2}$ (0.37) 39%
Quo.Vadis 0.64 (0.73) 1.5 ×$\times 10^{-2}$ (0.25) 15%
SLIFER No Dyn. 0.94 (0.95) 2.5 ×$\times 10^{-2}$ (0.14) 5.3%
SLIFER 0.94 (0.97) 4.1 ×$\times 10^{-2}$ (0.38) 19%
@@
Module TPR FPR ER
Signatures 0.34 $8.9\times 10^{-3}$ 0
MalConv 0.74 $4.8\times 10^{-2}$ 0
GBDT 0.75 (0.75) 2.8 (3.6) ×$\times 10^{-3}$ 0.036%
Nebula 0.41 (0.86) 7.8 ×$\times 10^{-3}$ (0.21) 45%
Quo.Vadis 0.80 (0.84) 9.2 ×$\times 10^{-3}$ (8.0 ×$\times 10^{-2}$) 4.8%
SLIFER No Dyn. 0.88 (0.88) 5.8 (5.9) ×$\times 10^{-2}$ 0.029%
SLIFER 0.89 (0.96) 6.4 ×$\times 10^{-2}$ (0.26) 20%
Table 3. Single modules detection results. Results in brackets are calculated considering detection errors as malware. We report the True Positive Rate (TPR), the False Positive Rate (FPR), and the Error Rate (ER).
@@
To better clarify the choice of the dynamic malware detector, we compare two state-of-the-art models: (i) Nebula, the transformer architecture described in Sect. 3.3; and Neurlux (Jindal et al., 2019). We plot the ROC curves of the two models, considering only successfully classified samples, dropping samples that resulted in classification errors. We see that Nebula outperforms Neurlux at each FPR, as shown in Fig. 3, that is the reason for our choice in the development of the dynamic module. Also, we want to calibrate our models finding a threshold at 1% FPR if possible, but Neurlux achieves very-high FPR at that threshold, which is unacceptable in a production detector. Thus, coherently with previous work (Trizna et al., 2023), we discarded Neurlux as possible module for SLIFER.
@@
Figure 3. ROC curve of Nebula and Neurlux, tested on $\mathcal{D}_{2}$. To compute these curves we discard impossible-to-analyse samples.
@@
As anticipated, we also include in the comparison the hybrid model Quo.Vadis (Trizna, 2022), considering that its approach of static and dynamic feature fusion is widespread in malware detection research. We compute metrics for Quo.Vadis considering a 1% FPR threshold computed on $\mathcal{D}_{2}$.
@@
Before drawing a conclusions on the predictive capabilities of tested models, we note that the best in terms of TPR and FPR are achieved by considering pre-processing errors as benignware. Thus, we are able to properly answer RQ1 for single models, by counter-intuitively asserting that, due to the majority of impossible-to-analyse samples among benignware, it is better to slightly decrease TPR to favour a low FPR regime. By looking at the best trade-off between FPR and TPR, we can not isolate a clear winner among models evaluated singularly. In particular, both GBDT-EMBER and Quo.Vadis achieve the best performance on both datasets, with extremely low false positive rates. It is worth to notice that Quo.Vadis also incorporates the same feature set provided by Anderson et al. (Anderson and Roth, 2018), which is also used for GBDT-EMBER. However, if we also include the error rate (ER) as part of our analysis, we can observe that GBDT-EMBER is clearly the winner of this analysis, with a very low percentage of crashes on both datasets. On the contrary, Quo.Vadis achieves a peak of 15% of pre-processing errors due to the combination of both static feature extraction and emulation. As competing candidate to Quo.Vadis, also MalConv shows matching performance with these models, by also being characterized with zero pre-processing errors thanks to it end-to-end structure.
@@
The best technique in terms of FPR is pattern-patching samples with YARA rules, but, as expected, the TPRs are very low for both datasets. In numbers, our sample dataset of rules comprising 2.7k signatures is only able to detect 750 malware in $\mathcal{D}_{1}$ and 6k in $\mathcal{D}_{2}$, scoring the least predictive detector. While this analysis highlights that, on average, each rule detects roughly 22 samples of $\mathcal{D}_{2}$, it is clear that YARA rules might match other models’ predictive capabilities only with a larger set of signatures. Lastly, dynamic analysis with emulation has a non-negligible number of errors, which hinders its already low metrics and its real impact in a multi-stage detector. This finding might be correlated with the fact that modeling the behavior of malware in an end-to-end perspective is a daunting task, thus confirming the superiority of static analysis (Dambra et al., 2023).
@@
5.2. SLIFER Classification
We now analyse the performances of SLIFER, with and without (SLIFER No Dyn.) the dynamic module as last step for the detection pipeline. This comparison will grant us insights on how much dynamic analysis improve the quality of predictions when used as a sequential detection mechanism, and we show the results in Table 3. For both settings compared to single modules, we clearly notice a drastic improvement in the trade-off between TPR and FPR for both datasets. While it is true that GBDT-EMBER achieves lower FPR, it is worth noticing that SLIFER increases TPR by 3% and 13% on $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ respectively. This increment is given by the sequential order of techniques of our methodology, that patches blind spots of single strategies thanks to the multitude of increasingly-accurate models. Thus, ER is lower compared to the dynamic module alone in both scenarios for two main reasons: (i) fewer malware samples that caused emulation to crash proceed to the last step of the pipeline, since they are stopped by earlier modules, and (ii) most of the pre-processing crashes are caused by benignware. Thus, we complete the answer to RQ1, by stating that also for SLIFER it is better to treat impossible-to-analyse samples as benign, to avoid higher FPR (which is 26% in the case of SLIFER Dyn.).
Sample label Backdoor Downloader Grayw. Miner Ransomw. Roguew. Spyware Virus Worm

### v3

@@ 3.4 SLIFER Implementation
We now present how we serialize the components we have described. SLIFER processes input programs by sequentially passing them from module to module, halting computations if one of those modules raises an alert. If one of the modules crashes due to pre-processing errors, SLIFER skips that module by passing the input to the next one. This can happen for EMBER-GBDT and Nebula, which require heavy pre-processing. We discuss classification error management in Sect. 5, by showing that it is better to label samples as benign in case of pre-processing crashes.
@@
For the machine learning static analysis module, we leverage a library called secml_malware (Demetrio and Biggio, 2021) [^8] https://github.com/pralab/secml_malware which wraps both MalConv and GBDT-EMBER models. Instead of training both from scratch, we leverage pre-trained open-source implementations of both  [^9] https://github.com/endgameinc/malware_evasion_competition. As the last module of SLIFER leverages the Nebula model provided by its original repository [^10] https://github.com/dtrizna/nebula. Emulation, which is the core of Nebula, is achieved through Speakeasy (Mandiant. Speakeasy: portable, modular, binary emulator designed to emulate Windows kernel and user mode malware, ), a Windows kernel emulation library, which generates a behavioral report in a JSON file.
4 Experimental Setting
@@
Datasets. In our experiments, we employed two different datasets, that deferred in cardinality, malware families, benignware, and period of collections. The first one ($\mathcal{D}_{1}\,$) is composed of 5k malware, collected from VirusTotal before year 2018, and 3.5k benignware, harvested from GitHub and clean Windows installations. We show its composition in Table 1, drafted again utilizing VirusTotal. As regards malware families, this dataset is unbalanced, with predominance of two specific families among all the others. While this composition might seem unsuitable for experiments, we believe it can be treated as a realistic snapshot of real production environments, where there is no control over the incoming threats to analyze.
Table 1: Composition of $\mathcal{D}_{1}\,$ divided by malware class. There are 551 samples we were unable to correctly label.
@@
Table 2: Composition of $\mathcal{D}_{2}\,$ Trizna (2022) by malware class.
@@
The second dataset ($\mathcal{D}_{2}\,$) reflects the test set of dynamic analysis records published by (Trizna, 2022), comprising 10K benign-ware and 17.5K malware samples spanned across seven types (such as ransomware, trojans, keyloggers, etc.) as shown in Table 2. The dataset was collected in April 2022 by partnering with an undisclosed security vendor. We re-collected PE files from public data sources based on released hashes to perform static and YARA level analysis applicable only to raw PE bytes. Contrary to $\mathcal{D}_{2}\,$, this dataset has perfect balances between malware families, making it a good baseline for fairly assessing performances across different techniques.
@@
False Positive Rate (FPR) to evaluate the rate of false alarms raised by each model/module. This is a key parameter in a production scenario, where high FPR compromises the usability of a malware detector.
F1-Score (F1) to evaluate whether SLIFER achieves high performances without excessive false positives or negatives.
Mean Detection Time (MDT) to evaluate the mean time that each component needs to analyze an input sample. We compute this metric on a subset of our datasets.
Error Rate (ER) of each component to estimate the percentage of impossible-to-analyze samples. We provide an experimental explanation of the best way to label those samples, as a decision must be taken when deployed.
@@
Comparison between SLIFER and single models. In the first experiment, we test the predictive capabilities of SLIFER on our two datasets, comparing it against each separate module. In doing this we motivate the choice of our proposed architecture, especially for the dynamic analysis module. We show that the chosen Nebula is better than a competitor model, named Neurlux (Jindal et al., 2019), a LSTM model trained on dynamic reports extracted from emulation. Also, we improve the comparison by including Quo.Vadis (Trizna, 2022) which is a hybrid-analysis model that merges both static and dynamic features. On the contrary to SLIFER, Quo.Vadis introduces a monolithic structure with single forward- and back-propagation path, processing static and dynamic components simultaneously, further employing a "meta-model" that accumulates representations from both analysis types to classify Windows executable samples.
@@
Effect of dynamic analysis in SLIFER Classification. Similarly to (Dambra et al., 2023), we investigate the efficacy of dynamic analysis. Different from their setting, here we use emulation as last step of the analysis without concatenating together both static and dynamic features. Thereafter, we test the performances of SLIFER in two cases: in the former, we do not include the dynamic module, and in the latter, we test the full pipeline, including the model chosen after the analysis of the two separate models. Lastly, we also include SLIFER performance on detection divided per malware family contained in $\mathcal{D}_{1}\,$ and $\mathcal{D}_{2}\,$.
Calibration of the components of SLIFER. We now introduce the technique we adopt to calibrate decision thresholds of the modules of SLIFER. We divide our two datasets $\mathcal{D}_{1}\,$and $\mathcal{D}_{2}\,$in two subsets, in which families are equally distributed: a validation set and a test set. In order to perform a fair calibration, we take half of the benignware samples from both $\mathcal{D}_{1}\,$and $\mathcal{D}_{2}\,$as well. We perform a simple version of an exhaustive grid search with this methodology: our hyperparameters in the grid are the three decision thresholds of the three ML modules in SLIFER (namely, MalConv, GBDT-EMBER and Nebula). For each combination, we calculate the F1-score on the validation set and take the threshold combination with the best F1-score. We repeat this process with different cardinalities of the validation set, aiming at picking the best strategy in terms of the F1-score and remaining test set dimension. We take the same research space for each decision threshold: =$\mathcal{R}=\{0.44,0.46,\dots,0.96,0.98\}$. Different from cross-validation, we do not re-train all the modules for each combination considering different folds, but we only evaluate scores on a validation set to properly tune decision thresholds. This methodology allows us to fairly compare the single modules and other state-of-the-art models with SLIFER, since we calibrate them too to the same FPR score as achieved by the "tuned" pipeline.
Effect of Calibration on SLIFER Classification. We perform the exhaustive grid search as described in the previous section, first halving our datasets into a validation and a test set. We compute the F1-score for each thresholds combination on the validation set and we repeat the process taking other percentages of the total amount of samples at our disposal. We analyze if calibration results significantly change increasing the number of samples used for calibration. After choosing the best split strategy and the best decision thresholds, we test the "calibrated" SLIFER on the remaining test set. We discuss obtained results comparing the "non-calibrated" pipeline performance to assess how much advantage we can obtain from a (simple) threshold-tuning procedure. For the sake of completeness, using the same validation set, we also select a threshold for every single model under test, picking the one that gives the same FPR as the calibrated SLIFER. With those thresholds, we repeat the same experiments with the single models.
@@
Detection Time Analysis. Here, we want to compare the mean speed of each separate module or model, aiming to analyze the advantages and disadvantages in terms of time. For the sake of completeness, we report the technical characteristics of the machine we used for this analysis: we use a workstation equipped with an Intel® Xeon(R) Gold 5420, two Nvidia L40 GPUs, and 540 GB of RAM. We calculate the MDT metric on two subsets of $\mathcal{D}_{1}\,$ and $\mathcal{D}_{2}\,$. The first one comprises 500 malware samples from $\mathcal{D}_{1}\,$ and 500 malware samples from $\mathcal{D}_{2}\,$. Similarly, we take 500 benignware samples from the two datasets. We also compute the standard deviation to show how the detection time can vary. For this evaluation, we separate malware and benignware samples because they differ in file dimension: benignware are usually bigger than malware, and this enlarges emulation, processing and classification times. Moreover, benign samples always pass through all SLIFER’s modules by design, and we prefer to evaluate times separately. To differ the two metrics, we denote with MDTm and MDTg the calculation of MDT for malware and benignware subsets respectively. Also, we compute detection times of SLIFER after calibration, since setting different thresholds impacts on where (by which module) a sample is detected.
@@
Pipeline Robustness Evaluation. We test SLIFER in terms of robustness, evaluating this metric with and without thresholds calibration. Due to the complexity of landing end-to-end attacks against pipelines like SLIFER, we opt for computing transfer attacks, by optimizing adversarial EXEmples (Demetrio et al., 2021b) – carefully-manipulated Windows programs aimed to evade detection – against a surrogate detector with similar capabilities, and then test them against SLIFER. Thus, we envision an attacker targeting the GBDT-EMBER model (which proves to be the best performing in most scenarios in terms of TPR and FPR, as we are going to discuss in Sect. 5), knowing is used inside the pipeline. To do so, we rely on GAMMA, a state-of-the-art black-box attack (Demetrio et al., 2021a) that injects non-executable portion of legitimate programs into malware. GAMMA leverages a genetic algorithm to select content that will included through either (i) the creation of new sections or (ii) appending bytes at the end (padding). GAMMA iteratively requests classification of perturbed samples to optimize the amount of injected content, until either evasion is achieved or the query budget is terminated.
Thus, we consider 1k samples classified as malicious by GBDT-EMBER from $\mathcal{D}_{2}\,$, and, as content to inject, we consider 75 benign sections picked from $\mathcal{D}_{2}\,$ benignware samples. Since GAMMA relies on regularization to control the total filesize of adversarial EXEmples, we set its parameter $\lambda=1\times 10^{-6}$, and we set the maximum number of queries to 100. These sections contain read-only data used by programs at runtime, but no further code is included inside adversarial EXEmples.
5 Experimental Results
We now describe all the findings we gathered from the experiments described in Sect. 4. To highlight the effect of different error-handling strategies, we will report in tables performance when flagging impossible-to-analyze samples both as malicious in brackets, and as benignware otherwise.
5.1 Single Modules Classification
@@
We test the performance of all the single modules with both $\mathcal{D}_{1}\,$ and $\mathcal{D}_{2}\,$ datasets in our possession, and we present results of each module in Table 3. We report the values of the described metrics when modules consider input as benignware in presence of pre-processing errors, by also including in brackets metrics computed considering classification errors as malware, as mentioned earlier in this section.
Table 3: Performance of single not-calibrated models and SLIFER. Results in brackets are calculated considering detection errors as malware, and benignware otherwise. We report the True Positive Rate (TPR), the False Positive Rate (FPR), F1 Score (F1), and the Error Rate (ER).
@@
Model TPR FPR F1 ER
Signatures 0.15 5.2 ×$\times 10^{-3}$ 0.27 0%
MalConv 0.78 2.0 ×$\times 10^{-2}$ 0.87 0%
GBDT 0.91 (0.92) 4.1 ×$\times 10^{-3}$ (0.12) 0.95 (0.91) 5.4%
Nebula 0.23 (0.61) 2.3 ×$\times 10^{-2}$ (0.37) 0.37 (0.64) 39%
Quo.Vadis 0.64 (0.73) 1.5 ×$\times 10^{-2}$ (0.25) 0.78 (0.77) 15%
SLIFER No Dyn. 0.94 (0.95) 2.5 ×$\times 10^{-2}$ (0.14) 0.96 (0.92) 5.3%
SLIFER 0.94 (0.97) 4.1 ×$\times 10^{-2}$ (0.38) 0.95 (0.87) 19%
@@
Model TPR FPR F1 ER
Signatures 0.34 $8.9\times 10^{-3}$ 0.51 0%
MalConv 0.74 $4.8\times 10^{-2}$ 0.84 0%
GBDT 0.75 (0.75) 2.8 (3.6) ×$\times 10^{-3}$ 0.85 (0.85) 0.036%
Nebula 0.41 (0.86) 7.8 ×$\times 10^{-3}$ (0.21) 0.58 (0.87) 45%
Quo.Vadis 0.80 (0.84) 9.2 ×$\times 10^{-3}$ (8.0 ×$\times 10^{-2}$) 0.89 (0.89) 4.8%
SLIFER No Dyn. 0.88 (0.88) 5.8 (5.9) ×$\times 10^{-2}$ 0.91 (0.91) 0.029%
SLIFER 0.89 (0.96) 6.4 ×$\times 10^{-2}$ (0.26) 0.93 (0.91) 20%
@@
To better clarify the choice of the dynamic malware detector, we compare two state-of-the-art models: (i) Nebula, the transformer architecture described in Sect. 3.3; and Neurlux (Jindal et al., 2019). We plot the ROC curves of the two models, considering only successfully classified samples belonging to $\mathcal{D}_{2}\,$, dropping the ones that resulted in classification errors. We see that Nebula outperforms Neurlux at each FPR, as shown in Fig. 3, that is the reason for our choice in the development of the dynamic module. Also, we want to calibrate our models finding a threshold at 1% FPR if possible, but Neurlux achieves very-low TPR at that threshold, which is unacceptable in a production detector. Thus, coherently with previous work (Trizna et al., 2024), we discarded Neurlux as possible module for SLIFER.
@@
Figure 3: ROC curve of Nebula and Neurlux, tested on $\mathcal{D}_{2}\,$. To compute these curves we discard impossible-to-analyze samples.
@@
As anticipated, we also include in the comparison the hybrid model Quo.Vadis (Trizna, 2022), considering that its approach of static and dynamic feature fusion is widespread in malware detection research. We compute metrics for Quo.Vadis considering a 1% FPR threshold computed on $\mathcal{D}_{2}\,$. We specify that stock MalConv and GBDT-EMBER models have a decision threshold computed at 1% FPR on their original test sets, while for Nebula we take the 1% FPR threshold from the ROC in Fig. 3.
Before drawing a conclusions on the predictive capabilities of tested models, we note that the best in terms of TPR, FPR, and F1 are achieved by considering pre-processing errors as benignware. Thus, we are able to properly answer RQ1 for single models, by counter intuitively asserting that, due to the majority of impossible-to-analyze samples among benignware, it is better to slightly decrease TPR to favour a low FPR regime.
@@
By looking at the best trade-off between FPR and TPR, we can not isolate a clear winner among models evaluated singularly. In particular, both GBDT-EMBER and Quo.Vadis achieve the best performance on both datasets, with extremely low false positive rates and good TPR. It is worth to notice that Quo.Vadis also incorporates the same feature set provided by (Anderson and Roth, 2018), which is also used for GBDT-EMBER. However, if we also include the error rate (ER) as part of our analysis, we can observe that GBDT-EMBER is clearly the winner of this analysis, with a very low percentage of crashes on both datasets. On the contrary, Quo.Vadis achieves a peak of 15% of pre-processing errors due to the combination of both static feature extraction and emulation. As competing candidate to Quo.Vadis, also MalConv shows matching performance with these models, by also being characterized with zero pre-processing errors thanks to it end-to-end structure.
@@
The best technique in terms of FPR is pattern-patching samples with YARA rules, but, as expected, the TPRs are very low for both datasets. In numbers, our sample dataset of rules comprising 2.7k signatures is only able to detect 750 malware in $\mathcal{D}_{1}\,$ and 6k in $\mathcal{D}_{2}\,$, scoring the least predictive detector. While this analysis highlights that, on average, each rule detects roughly 22 samples of $\mathcal{D}_{2}\,$, it is clear that YARA rules might match other models’ predictive capabilities only with a larger set of signatures. Lastly, dynamic analysis with emulation has a non-negligible number of errors, which hinders its already low metrics and its real impact in a multi-stage detector, while also inflating the number of false positives or negative depending on the error-handling strategy we consider. This finding might be correlated with the fact that modeling the behavior of malware in an end-to-end perspective is a daunting task, thus confirming the superiority of static analysis (Dambra et al., 2023).
@@
5.2 SLIFER Classification Performance
We now analyze the performances of SLIFER, with and without (SLIFER No Dyn.) the dynamic module as last step for the detection pipeline. This comparison will grant us insights on how much dynamic analysis improve the quality of predictions when used as a sequential detection mechanism, and we show the results in Table 3. For both settings compared to single modules, we clearly notice a drastic improvement in the trade-off between TPR and FPR for both datasets. While it is true that GBDT-EMBER achieves lower FPR, it is worth noticing that SLIFER increases TPR by 3% and 13% on $\mathcal{D}_{1}\,$ and $\mathcal{D}_{2}\,$ respectively. This increment is given by the sequential order of techniques of our methodology, that patches blind spots of single strategies thanks to the multitude of increasingly-accurate models. Thus, ER is lower compared to the dynamic module alone in both scenarios for two main reasons: (i) fewer malware samples that caused emulation to crash proceed to the last step of the pipeline, since they are stopped by earlier modules, and (ii) most of the pre-processing crashes are caused by benignware. Thus, we complete the answer to RQ1, by stating that also for SLIFER it is better to treat impossible-to-analyze samples as benign, to avoid higher FPR (which is 26% in the case of SLIFER Dyn.).
Table 4: True Positive Rate (TPR) of non-calibrated SLIFER divided by malware family inside in $\mathcal{D}_{1}\,$. Results in brackets are calculated considering detection errors as malware. They are considered benignware otherwise.
Architecture Backdoor Downloader Grayw. Miner Ransomw. Roguew. Spyware Virus Worm